OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 55°, a = 12, b = 14 A)B = 72.9°, C = 52.1°, c = 12.5; B = 107.1°, C = 17.9°, c = 12.5 B)B = 72.9°, C = 52.1°, c = 11.6; B = 107.1°, C = 17.9°, c = 4.5 C)B = 17.1°, C = 107.9°, c = 10.3; B = 162.9°, C = 72.1°, c = 10.3 D) B = 17.1°, C = 107.9°, c = 13.9; B = 162.9°, C = 72.1°, c = 13.9

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin(A)}{a}= \frac{\sin(B)}{b} ~~~~⇒~~~~ \frac{\sin(55)}{12}= \frac{\sin(B)}{14} }\)

OpenStudy (magbak):

i hope this helps http://www.mathportal.org/calculators.php